OpenStudy (mathmath333):

There is a natural number that becomes equal to the square of a natural number when 100 is added to it and to the square of another natural number when 169 is added to it .Find the number?

OpenStudy (anonymous):

x+100=n^2 x+169 = m^2

OpenStudy (cwrw238):

x + 100 = y^2 x + 169 = z^2

OpenStudy (anonymous):

diphontane equation

OpenStudy (cwrw238):

subtract z^2 - y^2 = 69 (z + y)(z - y) = 69 can you continue?

OpenStudy (cwrw238):

x y and z are all natural numbers

OpenStudy (mathmath333):

13,10.....?

OpenStudy (mathmath333):

23.3

OpenStudy (cwrw238):

z + y =23 z - y = 3 is a possibility

OpenStudy (mathmath333):

do i have to guess the answer?

OpenStudy (cwrw238):

no try and fit those numbers to the 2 equations

OpenStudy (cwrw238):

solve the 2 simultaneous equations first

OpenStudy (mathmath333):

z=13 y=10

OpenStudy (cwrw238):

yep

OpenStudy (cwrw238):

now work out value of x

OpenStudy (mathmath333):

how?

OpenStudy (cwrw238):

from the original equations

OpenStudy (cwrw238):

the other possibility is y + z= 69 y - z = 1

OpenStudy (cwrw238):

x + 100 = 10^2 x = 0 now to be honest I cant remember if 0 is a natural number

OpenStudy (mathmath333):

y=35 z=34

OpenStudy (cwrw238):

you can do the same processing for z + y = 69 z - y = 1 this gives z = 35 y = 34

OpenStudy (mathmath333):

925..?

OpenStudy (cwrw238):

giving x = 34^2 - 100

OpenStudy (cwrw238):

1056

OpenStudy (cwrw238):

sorry one of my last posts was incorrect its z - y = 1 not y - z = 1

OpenStudy (cwrw238):

there are 2 possibilities for (z + y)(z - y) = 69 if z and y are natural numbers the left hand side is either 23*3 or 69 * 1